# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

### v1

@@ 3.3.1 SiPM Effect Modeling
Gain 3000 ADC Count to 5000 ADC Count 3000 ADC Cout
Charge Resolution  15% 15%
Time Offset Unknown 0 ns or 20 ns
Time Spread  1 ns 1 ns
PDE 45% to 55% 50%
Internal Crosstalk < 1% 15%
External Crosstalk 20% to 30% 24.47%
Afterpulse < 1% 1%
3.3.2 Event Triggering
After converting the SiPM hits into electronic pulses and adding the pulses from SiPM effect, we can perform event triggering. The TAO electronics support two trigger modes: the nhit self-trigger and the external trigger. The nhit self-trigger scans a 300 ns time window along the time axis and triggers when the number of pulses within this window exceeds a threshold of 1000; the left boundary of the window is designated as the trigger time. The external trigger is forced to trigger at the start time of the detector simulation event.
3.3.3 Waveform Simulation
After triggering, the pulses from the trigger event are converted into waveforms. Each pulse is converted into a SPE waveform with the waveform starting point determined by the the pulse’s timestamp, and the waveform amplitude corresponding to the pulse’s amplitude. Subsequently, white noise is added with an amplitude equal to 10% of the SPE waveform amplitude.
After generating the waveforms, the charge and time of each waveform can be calculated. When three consecutive points exceed the threshold (set at 50% of the SPE amplitude), the waveform time is determined by the intersection of the line (fitted to the rising edge of the waveform) and the baseline. The charge is obtained by integrating the waveform backward from the threshold-crossing point to the baseline and forward to the baseline.
3.4 Simulated Data Sets
The simulated data sets used for SiPM calibration in this study are listed in Table 2. The DCR, IOCT and gain can be calibrated using arbitrary event types. In this study, we use electrons uniformly distributed within the LS with energies following a uniform distribution in the range of 1–10 MeV to represent arbitrary events. For the time offset calibration, the calibration source must be located at the center of the LS. This placement ensures that photons arrive simultaneously at all the SiPMs, thereby eliminating systematic bias induced by photon flight time. Calibration of the relative PDE requires that the incident photon flux on each SiPM is statistically uniform. Therefore, the calibration source is deployed at the center of the LS to ensure that the light field is isotropic. The $^{68}{\rm Ge}$ source fulfills this criterion, as it produces oppositely directed $\gamma$-ray pairs from positron annihilation, and the light emission from the LS is isotropic. For the calibration of EOCT, the SiPM-switching method is adopted, which requires the use of LED-based external triggering. The specifics of this methodology are elaborated in Section 6.
In the simulated data, the true values of all parameters are known. Accordingly, calibration accuracy is quantified as the bias and standard deviation. The calibrated bias is calculated as the average difference between the calibrated and true simulated values across all channels, and the calibrated standard deviation is obtained through the standard deviation of each channel calibrated value after subtracting the corresponding simulated true value (excluding the bias of relative PDE).
Parameter Particle Position Trigger Type Data Type True Value
[…]
Gain Arbitrary Arbitrary Self Trigger Charge 3000 ADC Count
Time Offset LED LS Centre External Trigger Hit Time 0ns or 20ns
Relative PDE $^{68}{\rm Ge}$ LS Centre Self Trigger Hit Time 1
Internal Crosstalk Arbitrary Arbitrary Self Trigger Charge 0.15
External Crosstalk LED LS Centre External Trigger Charge From EOCT disabled data
@@ 4.1 Dark Count Rate
[…]
The DCR quantifies the level of DN of SiPM and is defined as the number of DN generated per unit sensitive area of the SiPM per unit time. It can be estimated by counting the hits occurring in the plateau before the peak of the main hits generated by the physics events, as shown in Figure 2, which displays the simulated hit time distribution of a channel. With $n$ denoted as the total number of the counted hits in the plateau (indicated by the red line in the figure), $N$ as the number of physics events, $t$ as the duration of the plateau before the trigger timestamp (from -100 ns to 0 ns), and $S$ as the sensitive area of the SiPM. Then, the DCR can be calculated using Equation 4.1:
$DCR=\frac{n}{N\cdot t\cdot S}$ (4.1)
Figure 2: The Simulated Hit Time Distribution for a Channel. The trigger algorithm slides a 300 ns time window along the time axis. An event is triggered when the number of channels with hit within this trigger time window exceeds a preset threshold (set to 1000 in the TAO electronics simulation). The trigger time corresponds to the start of the trigger time window and is aligned to t = 0 ns. The scintillation light from the physics event is emitted within a short time after the trigger time, so hits occurring before the trigger time primarily originate from DN.
However, the hits observed in the plateau range may also include minor contributions from hits generated by the EOCT of DN from other SiPMs, as well as hits from delayed APs. For the TAO SiPM, the delayed AP has been measured to be very weak [Ref26]. Therefore, only EOCT contamination needs to be considered in the DCR calibration.
@@
[…]
@@ 4.2 Time Offset
The hit time of the SiPM obtained via the electronic readout satisfies Equation 4.2, where $t_{R}$ is the hit time from the electronic readout. It is calculated as $t_{R}=t_{A}-t_{T}$, where $t_{A}$ is the absolute timestamp and $t_{T}$ is the trigger timestamp. The absolute timestamp consists of several components: $t_{P}$ is the time when the particle reaches and deposits energy in the LS; $t_{\rm LS}$ is the time when the LS emits scintillation photons; $t_{\rm tof}$ is the time of flight of scintillation photons to the SiPM; the electronic processing time, which is divided into two parts: $t_{\rm same}$ and $t_{\rm timeoffset}$. Here, $t_{\rm same}$ represents the identical portion of the processing time across all electronic channels, while the differing portions constitute the time offset ($t_{\rm timeoffset}$) in different channels.
@@
Since the TAO experiment only requires correcting the relative time offsets between electronic channels [Ref15] to ensure that the timestamps of all channels are synchronized, this study deploys an LED source at the center of the LS to calibrate the relative time offset between different channels. The LED source acquires events using an external trigger method, ensuring that the trigger time $t_{T}$ is identical across all channels. Therefore, $t_{T}$ does not affect the calibration of relative time offset. The $t_{P}$ and $t_{\rm LS}$ are statistically identical across all channels, so the $t_{P}$ and $t_{\rm LS}$ does not influence relative time offset calibration. Because the LED is deployed at the center of the LS, the distance from the LS center to each SiPM channel is the same, meaning the $t_{\rm tof}$ is identical across all channels. Consequently, the relative time offsets between different channels can be derived from the relative $t_{R}$ differences for each channel.
@@
Figure 4: Double-Sided Crystal Ball Function Fit to the Time Spectrum.
@@
Figure 5: Time Offset Calibration Result. The left figure shows the calibration result for each channel. In this simulation, half of the channels are set to the time offset of 0 ns, while the other half are set to the time offset of 20 ns. The middle and right figures show the bias between the time offset calibration result and the mock truth.
In the actual TAO CD, the the calibration source’s position may be offset from its deployed position, and this offset contributes to a bias in the time of flight. Specifically, a 1 cm offset results in a 0.13 ns bias between the SiPMs located at the south and north poles of the CD. Considering that the time resolution of the TAO electronics exceeds 1 ns [Ref15], both the calibration bias and the calibration source’s positional offset from the LS center are acceptable.
@@ 4.3 Relative Photon Detection Efficiency
As shown in Equation 4.4, the photon detection efficiency (PDE) of a SiPM is defined as the ratio of the number of photons detected by the SiPM ($N_{\rm Det}$) to the number of photons incident on the SiPM surface ($N_{\text{photon}}$). The PDE is determined by the SiPM’s quantum efficiency (QE), electron avalanche probability, and geometric fill factor [Ref27, Ref28]. Specifically, the QE depends on the wavelength of the photon, the electron avalanche probability is a function of the operating voltage of the SiPM, and the geometric fill factor is constant for each SiPM.
$PDE=\frac{N_{\rm Det}}{N_{\text{photon}}}$ (4.4)
Since the number of photons incident on the SiPM surface depends on the particle type, particle energy, light yield of the LS, and light attenuation length of the LS, calibrating the absolute PDE of each SiPM is impractical. However, calibrating the relative PDE is feasible. The TAO calibration system employs three types of calibration sources: an LED source, a $^{68}{\rm Ge}$ source and a combined $\gamma$ source [Ref21]. The light field produced by LED source may be not isotropic, and the combined $\gamma$ source consists of multiple $\gamma$ sources with different energies. In contrast, the $^{68}{\rm Ge}$ source is a positron annihilation $\gamma$ source. Due to the presence of the source shell, positrons annihilate inside the shell, producing a pair of $\gamma$ emitted in opposite directions, which enhances the uniformity of the light field. Therefore, the $^{68}{\rm Ge}$ source is used for relative PDE calibration.
Deploying the $^{68}{\rm Ge}$ source at the center of the LS, the number of photons detected by each channel follows a Poisson distribution. The parameter of the Poisson distribution is $N_{\rm Det}$ which represents the mean number of detected photons. It satisfies Equation 4.2 where $N_{\text{PE}=0}$ is the number of events with zero hits in the channel, and $N_{\text{Event}}$ is the number of events. Since the $^{68}{\rm Ge}$ source is deployed at the center of the LS and the light emission from the LS is isotropic, the average number of incident photons on the surface of each channel is the same. According to Equation 4.3, the relative PDE of each channel corresponds to the relative $N_{Det}$ value.
@@
However, the DN effect can also generate hits. If the DCR of different channels vary, the values of $N_{\text{PE}=0}$ will differ, causing biases in the relative PDE calculated based on the relative $N_{\rm Det}$. Therefore, it is necessary to correct $N_{\text{PE}=0}$ before calculating the relative $N_{\rm Det}$. As shown in Figure 2, after correcting for the time offset effect, we count the hits in the time range from 0 ns to 160 ns, considering them as contributions from DN, and denote the number of events with hits in this DN time range as $N_{\rm DN}$. We also count the hits in the time range from 240 ns to 400 ns, considering them as contributions from both DN and $^{68}{\rm Ge}$, and denote the number of events with hits in the $^{68}{\rm Ge}$ time range as $N_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$. As shown in Equation 4.6, $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$ is the probability of hits generated by DN, and $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$ is the probability of hits generated by both DN and $^{68}{\rm Ge}$. Since DN and $^{68}{\rm Ge}$ are independent, the combined probability is the sum of the probabilities of hits generated by DN alone, by $^{68}{\rm Ge}$ alone, and by both simultaneously. Therefore, the probability of hits generated by Ge68 $P_{\rm{}^{68}{\rm Ge}}$ can be derived from $P_{\rm DN}$ and $P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}$. In this simulation, we set four different DCR values: 20, 40, 60, and 80 ${\mathrm{mm}}^{2}$. The results before and after applying this correction method are shown in Figure 6, this method can effectively correct the bias caused by different DCR levels.
$\begin{split}P_{\rm DN}&=\frac{N_{\rm DN}}{N_{\text{Event}}}\\ P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}&=\frac{N_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}}{N_{\text{Event}}}\\ &=P_{\rm DN}\cdot(1-P_{\rm{}^{68}{\rm Ge}})+P_{\rm{}^{68}{\rm Ge}}\cdot(1-P_{\rm DN})+P_{\rm{}^{68}{\rm Ge}}\cdot P_{\rm DN}\\ P_{\rm{}^{68}{\rm Ge}}&=\frac{P_{\rm{}^{68}{\rm Ge}_{\rm WithDN}}-P_{\rm DN}}{1-P_{\rm DN}}\\ N_{\text{PE}=0}&=(1-P_{\rm{}^{68}{\rm Ge}})\cdot N_{\text{Event}}\end{split}$ (4.6)
Figure 6: Calibration Bias Before and After DCR Correction. The left figure shows the bias which uncorrected DN differences, and the right figure shows bias which corrected.
@@
After Correcting the DN differences in each channel, the relative PDE calibration result is shown in the left part of Figure 7. There is an approximate 3% bias between the channels at the pole and equator, which is caused by optical reflection on the SiPM surface. As shown in the right part of Figure 7, this bias disappears after disabling the optical reflection on the SiPM surface in the detector simulation. This bias will be corrected based on the result of the detector optical model and simulation in the future.
Figure 7: Bias Between Relative PDE Calibration Result and Mock Truth. The left figure shows the results under normal configuration, while the right figure shows those with the optical reflection effect on SiPM surface disabled.

### v2

@@ 3.4.5 Optical Crosstalk
where $k$ represents the total number of SiPM electrical pulses induced by one primary avalanche signal and all subsequent OCT events triggered by the primary signal, and $\lambda_{\rm OCT}$ is the Poisson parameter characterizing the intrinsic OCT probability per avalanche event.
The physical meaning behind this mathematical formulation can be intuitively interpreted as a multi-generation cascaded avalanche process: a primary avalanche signal produces a set of first-generation OCT events through a Poisson process with rate parameter $\lambda_{\rm OCT}$. Each of these first-generation OCT events will trigger an independent avalanche, which can further generate second-generation OCT events via an identical Poisson process with the same rate parameter $\lambda_{\rm OCT}$. This iterative cascade process continues until no additional OCT events are generated in a given generation. Mathematically, the total count of the primary signal and all cascaded OCT events across all generations converges to the Borel distribution given above.
More generally, for avalanche signals generated by a Poisson process with a Poisson coefficient of $\mu$, the sum of the number of avalanche signals and the number of OCT events generated by the avalanche signals mathematically follows a generalized Poisson distribution:
In the TAO electronics simulation, IOCT and EOCT are simulated separately. For IOCT, the corresponding number of IOCT pulses is generated via Borel distribution sampling. IOCT will be induced by every electrical pulse generated from a SiPM photon hit, a dark noise electrical pulse, and an EOCT electrical pulse. According to the study by Guan et al. [18], the IOCT of HPK SiPMs in linear alkylbenzene is very low, with a Poisson coefficient $\lambda_{\rm IOCT}$ less than 0.1%, which makes this effect negligible and no corresponding calibration required. The TAO SiPM batch tests measured the in-air IOCT probability of the SiPMs, which value is approximately 15%. We directly adopted this value in the TAO electronics simulation, to investigate how to calibrate this parameter under the scenario where the IOCT effect is non-negligible.
The simulation of EOCT is implemented by performing Poisson sampling on each generation of electrical pulses. Each electrical pulse, including those induced by SiPM photon hits, dark noise, IOCT, is subjected to Poisson sampling with a coefficient of $\lambda_{\rm EOCT}$ to generate first-generation EOCT pulses. These newly generated EOCT pulses, together with the IOCT produced by them, are further processed with Poisson sampling of the same coefficient $\lambda_{\rm EOCT}$ to produce a new generation of EOCT pulses. Such a cascaded process proceeds iteratively until no additional EOCT pulses are generated. In the TAO electronics simulation, the $\lambda_{\rm EOCT}$ is adopted from the measurement of EOCT for HPK SiPMs reported in the study by Guan et al. [18], with a value of 0.2447.
The electronics channel ID hit by EOCT is obtained by simulating a specific photon emission distribution on the SiPM surface and observing the photon hit responses of other SiPMs. Taking the SiPM with SiPM ID = 0 as an example, one million optical photons are simulated using a particle gun on the surface of the corresponding SiPM. The angle between the photon propagation direction and the surface normal is sampled from a Gaussian distribution with a mean of $0^{\circ}$ and a standard deviation of $30^{\circ}$. If the sampled angle larger than $90^{\circ}$, the sampling is repeated until angle less than or equal to $90^{\circ}$.
Following the above simulation procedure, the distribution of SiPM ID hit by EOCT generated from the SiPM with ID = 0 is shown in Figure 3. Since the photons generated by EOCT are emitted predominantly along the normal direction of the SiPM surface, the SiPM facing the EOCT-emitting SiPM within each layer of the SiPMs receive the largest number of EOCT photons. This results in a multi-peak structure in the SiPM ID distribution, where each peak corresponds to the EOCT response received by the SiPMs in each layer. The electronics channel ID hit by EOCT is then obtained by sampling from this histogram.
Figure 3: Hit SiPM ID Distribution from EOCT Hits Generated by the SiPM ID = 0 Channel.
It is worth noting that since the coverage of TAO SiPMs is not 100%, the total number of photon hits is less than one million. The ratio of the total number of photon hits to one million is defined as the geometric factor for EOCT. The actual Poisson coefficient for EOCT is 0.2447 multiplied by the geometric factor corresponding to the SiPM. The geometric factor varies for SiPMs at different positions. The distribution of the geometric factor for all 4024 SiPMs is shown in Figure 4, the mean value is 0.7429.
Figure 4: Geometric Factor Distribution for EOCT.
3.4.6 Afterpulse
During the avalanche discharge process of the APD, there is a probability that electron-hole pairs are trapped by lattice defects. The trapped carriers will be re-released, which in turn triggers a new avalanche process; this phenomenon is AP [35]. The AP in SiPMs can be divided into two categories. One is delayed AP, which has a long time interval from the main avalanche signal (> 1 $\mu\text{s}$). For the TAO SiPM, the measured delayed AP is negligible [45]. The other is non-delayed AP. Due to the short time interval between this type of AP and the main signal, the APD is still in the avalanche recovery phase at this time, so the gain of the AP is lower than that of the primary avalanche pulse.
Based on the result of the TAO SiPM mass test [42], the AP probability of the TAO SiPM is very low (< 1%). In the TAO simulation, we have implemented a 1% AP effect. The time interval $\Delta t$ between the AP and the primary pulse, as well as the relative gain of the AP $\text{Gain}_{AP}$ with respect to the primary pulse, both follow an exponential distribution [10], as shown in Equation 3.4 and Equation 3.5, respectively.
$f(\Delta t)=e^{-\frac{\Delta t}{\tau_{AP}}}$ (3.4)
$\text{Gain}_{AP}=1-e^{-\frac{\Delta t}{\tau_{rec}}}$ (3.5)
where $\tau_{\text{AP}}$ is the AP time constant, and $\tau_{\text{rec}}$ is the avalanche recovery time constant. Since these two time constants were not measured in the TAO SiPM mass test, we adopted the values reported in Ref. [34], which are 20 ns and 13 ns, respectively.
@@ 4.1 Dark Count Rate
The DCR can be estimated by counting the recorded hits in the plateau region before the main signal hits from physics events. This plateau region corresponds to the pre-trigger time, which is the area to the left of the red line (indicating the trigger time) in Figure 5. The DCR can be calculated using Equation 4.1:
$\text{DCR}=\frac{n}{N\cdot t\cdot S}$ (4.1)
Where $n$ denoted as the total number of the counted hits in the plateau, $N$ as the number of physics events, $t$ as the duration of the plateau before the trigger timestamp (from -100 ns to 0 ns), and $S$ as the sensitive area of the SiPM.
Figure 5: Hit time distribution from self-triggered events. The trigger time corresponds to 0 ns, which is indicated by the red line in the figure. The hits from physics events are concentrated around 300 ns, as determined by the 300 ns width of the trigger time window. The plateau preceding the hits of physics events corresponds to the dark noise plateau.
However, the hits observed in the plateau range may also include minor contributions from hits generated by the EOCT of dark noise from other SiPMs, as well as hits from delayed APs. For the TAO SiPM, the delayed AP has been measured to be very weak [45]. Therefore, only EOCT contamination needs to be considered in the DCR calibration.
@@
Figure 6: DCR Calibration Results. The left two panels: calibration results with EOCT included. The right two panels: calibration results with EOCT excluded. The upper two panels: comparison between the calibration results and the simulation truth in each channel. The lower two panels: 1D histogram filled with each bin value from the upper 2D histogram.
@@ 4.2 Time Offset
The hit time of the SiPM obtained via the electronic readout satisfies Equation 4.2, where $t_{R}$ is the hit time from the electronic readout. It is calculated as $t_{R}=t_{A}-t_{T}$, where $t_{A}$ is the absolute timestamp and $t_{T}$ is the trigger timestamp. The absolute timestamp consists of several components: $t_{P}$ is the time when the particle reaches and deposits energy in the liquid scintillator; $t_{\rm LS}$ is the time when the liquid scintillator emits scintillation photons; $t_{\rm tof}$ is the time of flight of scintillation photons to the SiPM; the electronic processing time, which is divided into two parts: $t_{\rm same}$ and $t_{\rm timeoffset}$. Here, $t_{\rm same}$ represents the identical portion of the processing time across all electronic channels, while the differing portions constitute the time offset ($t_{\rm timeoffset}$) in different channels.
@@
Since the TAO experiment only requires correcting the relative time offsets between electronic channels [2] to ensure that the timestamps of all channels are synchronized, this study deploys an LED source at the center of the liquid scintillator to calibrate the relative time offset between different channels. The LED source acquires events using an external trigger method, ensuring that the trigger time $t_{T}$ is identical across all channels. Therefore, $t_{T}$ does not affect the calibration of relative time offset. The $t_{P}$ and $t_{\rm LS}$ are statistically identical across all channels, so the $t_{P}$ and $t_{\rm LS}$ does not influence relative time offset calibration. Because the LED is deployed at the center of the liquid scintillator, the distance from the liquid scintillator center to each SiPM channel is the same, meaning the $t_{\rm tof}$ is identical across all channels. Consequently, the relative time offsets between different channels can be derived from the relative $t_{R}$ differences for each channel.
@@
Figure 7: Double-Sided Crystal Ball Function Fit to the Time Spectrum. The mean value of the fitted function corresponds to the $t_{R}$ value.
@@
Figure 8: Time Offset Calibration Results. The upper panel: channel-level calibration results. The middle panel: comparison between the calibration results and the simulation truth in each channel. Lower panel: 1D histogram filled with the each bin value from the middle 2D histogram.
In the actual TAO CD, the the calibration source’s position may be offset from its deployed position, and this offset contributes to a bias in the time of flight. Specifically, a 1 cm offset results in a 0.13 ns bias between the SiPMs located at the south and north poles of the CD. Considering that the time resolution of the TAO electronics exceeds 1 ns [2], both the calibration bias and the calibration source’s positional offset from the liquid scintillator center are acceptable.
@@ 4.3 Relative Photon Detection Efficiency
As shown in Equation 4.4, the PDE of a SiPM is defined as the ratio of the number of photons detected by the SiPM ($N_{\rm Det}$) to the number of photons incident on the SiPM surface ($N_{\text{photon}}$):
$\text{PDE}=\frac{N_{\rm Det}}{N_{\text{photon}}}$ (4.4)
Since the number of photons incident on the SiPM surface depends on the particle type, particle energy, light yield of the liquid scintillator, and light attenuation length of the liquid scintillator, calibrating the absolute PDE of each SiPM is impractical. However, calibrating the relative PDE is feasible. The TAO calibration system employs three types of calibration sources: an LED source, a $^{68}{\rm Ge}$ source and a combined $\gamma$ source [46]. The light field produced by LED source may be not isotropic, and the combined $\gamma$ source consists of multiple $\gamma$ sources with different energies. In contrast, the $^{68}{\rm Ge}$ source is a positron annihilation $\gamma$ source. Due to the presence of the source shell, positrons annihilate inside the shell, producing a pair of $\gamma$ emitted in opposite directions, which enhances the uniformity of the light field. Therefore, the $^{68}{\rm Ge}$ source is used for relative PDE calibration.
Deploying the $^{68}{\rm Ge}$ source at the center of the liquid scintillator, the number of photons detected by each channel follows a Poisson distribution. The mean number of photons detected by the SiPM is the parameter of the Poisson distribution, which satisfies:
@@
Where $N_{\text{PE}=0}$ is the number of events without $^{68}{\rm Ge}$ photon hits in the channel, and $N_{\text{Event}}$ is the number of events. Since the $^{68}{\rm Ge}$ source is deployed at the center of the liquid scintillator and the light emission from the liquid scintillator is isotropic, the average number of incident photons on the surface of each channel is the same. According to Equation 4.4, the relative PDE of each channel corresponds to the relative $N_{Det}$ value.
However, the dark noise effect can also generate hits. If the DCR varies across different readout channels, the number of hits contributed by dark noise will differ between channels, which in turn leads to variations in $N_{\text{PE}=0}$, thereby introducing a bias in the calibration results of the relative PDE. For this reason, a dedicated correction for the dark noise effect is required.
The hit time distribution is shown in Figure 5. After applying the time offset correction, we count hits from 0 ns to 160 ns, which are purely contributed by dark noise. Then, also count the hits from 240 ns to 400 ns, which include contributions from both dark noise and $^{68}{\rm Ge}$. The probability of a dark noise occurring within the 160 ns time window, denoted as $P_{\rm DN}$, is defined as the ratio of the number of events containing dark noise $N_{\rm DN}$ to the $N_{\text{Event}}$. The probability of observing either dark noise or $^{68}{\rm Ge}$ (or both), denoted as $P_{\rm Ge,DN}$, is defined as the ratio of the number of events containing dark noise, or $^{68}{\rm Ge}$, or both occurring $N_{\rm Ge,DN}$ to the $N_{\text{Event}}$.
[…]
Dark noise and $^{68}{\rm Ge}$ are statistically independent. Therefore, the probability of the union of these two events $P_{\rm Ge,DN}$ equals the sum of the probabilities of three mutually exclusive scenarios: dark noise occurs without $^{68}{\rm Ge}$, $^{68}{\rm Ge}$ occurs without dark noise, and both dark noise and $^{68}{\rm Ge}$ occur simultaneously.
$\begin{split}P_{\rm Ge,DN}=P_{\rm DN}\cdot(1-P_{\rm Ge})+P_{\rm Ge}\cdot(1-P_{\rm DN})+P_{\rm Ge}\cdot P_{\rm DN}\end{split}$ (4.7)
Meanwhile, the probability of no $^{68}{\rm Ge}$ occurring within the 240 ns to 400 ns time window, denoted as $1-P_{\rm Ge}$, is defined as the ratio of the $N_{\text{PE}=0}$ to the $N_{\text{Event}}$.
$\begin{split}N_{\text{PE}=0}&=(1-P_{\rm Ge})\cdot N_{\text{Event}}\end{split}$ (4.8)
where $P_{\rm Ge}$ is derived from Equation 4.6 and Equation 4.7, and is given by:
$\begin{split}P_{\rm Ge}&=\frac{P_{\rm Ge,DN}-P_{\rm DN}}{1-P_{\rm DN}}\end{split}$ (4.9)
In this simulation, we set four different DCR values: 20, 40, 60, and 80 ${\mathrm{mm}}^{2}$. The results before and after applying this correction method are shown in Figure 9, this method can effectively correct the bias caused by different DCR levels.
Figure 9: Relative PDE Calibration Results Before and After DCR Correction. The left and right panels present the comparison between the PDE calibration results and the simulation truth before and after DCR correction, respectively.
Figure 10: Relative PDE Calibration Results Before and After Accounting for the Optical Reflection Effect of the SiPM. The left two panels: calibration results with the SiPM optical reflection effect included. The right two panels: calibration results with the SiPM optical reflection effect excluded. The upper two panels: comparison between the calibration results and the simulation truth in each channel. The lower two panels: 1D histogram filled with each bin value from the upper 2D histogram.
@@
After Correcting the dark noise differences in each channel, the relative PDE calibration result is shown in the left part of Figure 10. There is an approximate 3% bias between the channels at the pole and equator, which is caused by optical reflection on the SiPM surface. As shown in the right part of Figure 10, this bias disappears after disabling the optical reflection on the SiPM surface in the detector simulation. This bias will be corrected based on the result of the detector optical model and simulation in the future.